Use the given information to evaluate each expression. (Enter your answers in exact form.)
cot s = 9 with π < s < 3π/2
a.) sin 2s
b.) cos 2s
c.) tan 2s

Answer by lwsshak3(11628)

Use the given information to evaluate each expression. (Enter your answers in exact form.)
cot s = 9 with π < s < 3π/2
Given angle s is in quadrant III, where sin<0, cos<0, tan>0
Hypotenuse of reference right triangle=√(1^2+9^2)=√(1+81)=√82
sin s=-1/√82=-√82/82
cos s=-9/√82=-9√82/82
tan s=1/9
..
a.) sin2s=2sin(s)cos(s)=2(-√82/82)(-9√82/82)=1476/6724
b.) cos2s=cos^2(s)-sin^2(s)=81/82-1/82=80/82
c.) tan2s=2tan(s)/(1-tan^2(s)=2/9/(1-1/81)=(2/9)/(80/81)=18/80